OpenStudy (anonymous):

What are the solutions of the equation? z² – 6z – 27 = 0

OpenStudy (karatechopper):

first u need to factor

OpenStudy (anonymous):

(x-9)(x+3)=0

OpenStudy (anonymous):

x=9, x=-3

OpenStudy (anonymous):

z^2 -9z+3z-27=0 z(z-9)+3(z-9)=0 (z+3)(z-9)=0 z=-3 or 9

OpenStudy (anonymous):

sorry its z